# Follow Your Instructions?

### v1

@@ 1 Introduction
Text-to-Image (T2I) generation has emerged as a cornerstone of multimodal AI, enabling the translation of abstract textual concepts into detailed visual content, advancing applications from digital art to scientific visualization. Recent T2I models can be categorized into two main paradigms. Diffusion-based methods—exemplified by Stable Diffusion [1, 2], PixArt [3, 4, 5], Playground [6, 7], FLUX [8], SANA [9, 10, 11], and others[12, 13, 14, 15]—leverage U-Net or Diffusion-Transformer backbones to iteratively denoise Gaussian noises into photorealistic images, delivering superior visual fidelity and diversity. On the other hand, autoregressive (AR) approaches such as LlamaGen [16], Janus [17, 18], Infinity/VAR [19, 20] and other influential open-source efforts [21, 22, 23] treat images as token sequences, employing next-token prediction or scale-progressive generation to synthesize images. Very recently, commercial T2I models such as GPT-4o [24], Imagen 3 [25] and MidJourney v7 [26] have propelled T2I to a new level. In particular, GPT-4o demonstrates powerful instruction-following capability, which can understand highly intricate prompts and return visually precise, stylistically coherent images in a single conversational loop.
@@ 4.1 Performance of T2I Models on TIIF-Bench
Model    Overall    Basic Following    Advanced Following    Designer          Avg    Attribute    Relation    Reasoning    Avg    Attribute +Relation    Attribute +Reasoning    Relation +Reasoning    Style    Text    Real World    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long Diffusion based Open-Source Models FLUX.1 dev    71.09    71.78    83.12    78.65    87.05    83.17    87.25    80.39    75.01    72.39    65.79    68.54    67.07    73.69    73.84    73.34    69.09    71.59    66.67    66.67    43.83    52.83    70.72    71.47 SD XL    54.96    42.13    65.72    53.28    59.33    50.83    77.57    62.57    60.32    46.57    49.73    36.22    47.82    35.57    56.22    45.34    52.59    36.09    73.33    60.00    16.83    0.83    50.92    41.59 SD 3    67.46    66.09    78.32    77.75    83.33    79.83    82.07    78.82    71.07    74.07    61.46    59.56    61.07    64.07    68.84    70.34    50.96    57.84    66.67    76.67    59.83    20.83    63.23    67.34 SD 3.5    71.15    66.96    78.34    79.56    79.50    76.50    80.96    83.21    72.46    78.71    67.67    61.18    66.46    61.89    73.53    74.15    60.03    61.53    73.33    63.33    70.52    42.52    64.43    66.39 SANA Sprint    63.68    58.50    76.58    71.00    75.33    71.33    81.82    72.07    72.57    69.57    57.67    51.80    55.32    54.94    68.46    66.72    62.59    63.46    80.00    60.00    8.82    5.83    66.96    58.01 SANA 1.5    67.15    65.73    79.66    77.08    79.83    77.83    85.57    83.57    73.57    69.82    61.50    60.67    65.32    56.57    69.96    73.09    62.96    65.84    80.00    80.00    17.83    15.83    71.07    68.83 Playground v2    45.64    52.78    59.83    69.58    51.33    66.33    70.57    76.07    57.57    66.32    38.43    44.75    41.57    45.57    48.96    59.97    41.72    54.84    53.33    60.00    0.00    0.83    45.32    46.44 Playground v2.5    47.73    54.82    63.08    68.08    57.83    73.83    71.82    77.32    59.57    53.07    40.73    48.17    39.70    45.82    49.59    64.22    44.22    46.72    60.00    80.00    0.00    4.83    47.19    47.56 PixArt-delta    41.01    48.24    53.83    59.25    46.33    52.83    62.07    71.32    53.07    53.57    34.60    42.77    32.44    37.44    53.59    56.59    36.96    49.46    46.67    73.33    0.00    0.00    38.23    40.10 PixArt-alpha    44.37    50.50    55.50    61.00    52.33    56.33    63.82    74.07    50.32    52.57    38.71    44.90    37.82    41.32    58.84    52.46    40.22    47.09    50.00    76.67    0.00    0.83    45.70    53.16 PixArt-sigma    62.00    58.12    70.66    75.25    69.33    78.83    75.07    77.32    67.57    69.57    57.65    49.50    65.20    56.57    66.96    61.72    66.59    54.59    83.33    70.00    1.83    1.83    62.11    52.41 LUMINA-Next    50.93    52.46    64.58    66.08    56.83    59.33    67.57    71.82    69.32    67.07    44.75    45.63    51.44    43.20    51.09    59.72    44.72    54.46    70.00    66.67    0.00    0.83    47.56    49.05 Hunyuan-DiT    51.38    53.28    69.33    69.00    65.83    69.83    78.07    73.82    64.07    63.32    42.62    45.45    50.20    41.57    59.22    61.84    47.84    51.09    56.67    73.33    0.00    0.83    40.10    44.20 AR based Open-Source Models Llamagen    41.67    38.22    53.00    50.00    48.33    42.33    59.57    60.32    51.07    47.32    35.89    32.61    38.82    31.57    40.84    47.22    49.59    46.22    46.67    33.33    0.00    0.00    39.73    35.62 LightGen    53.22    43.41    66.58    47.91    55.83    47.33    74.82    45.82    69.07    50.57    46.74    41.53    62.44    40.82    61.71    50.47    50.34    45.34    53.33    53.33    0.00    6.83    50.92    50.55 Show-o    59.72    58.86    73.08    75.83    74.83    79.83    78.82    78.32    65.57    69.32    53.67    50.38    60.95    56.82    68.59    68.96    66.46    56.22    63.33    66.67    3.83    2.83    55.02    50.92 Infinity    62.07    62.32    73.08    75.41    74.33    76.83    72.82    77.57    72.07    71.82    56.64    54.98    60.44    55.57    74.22    64.71    60.22    59.71    80.00    73.33    10.83    23.83    54.28    56.89 JanusPro    66.50    65.02    79.33    78.25    79.33    82.33    78.32    73.32    80.32    79.07    59.71    58.82    66.07    56.20    70.46    70.84    67.22    59.97    60.00    70.00    28.83    33.83    65.84    60.25 Closed-Source Models DALL-E 3    74.96    70.81    78.72    78.50    79.50    79.83    80.82    78.82    75.82    76.82    73.39    67.27    73.45    67.20    72.01    71.34    63.59    60.72    89.66    86.67    66.83    54.83    72.93    60.99 MidJourney v6    70.78    67.70    76.00    69.08    77.83    69.33    81.32    73.07    68.82    64.82    68.54    67.62    57.82    61.95    69.84    63.96    57.46    60.34    83.33    73.33    75.83    73.83    65.10    68.46 MidJourney v7    68.74    65.69    77.41    76.00    77.58    81.83    82.07    76.82    72.57    69.32    64.66    60.53    67.20    62.70    81.22    71.59    60.72    64.59    83.33    80.00    24.83    20.83    68.83    63.61 FLUX.1 Pro    67.32    69.89    79.08    78.91    78.83    81.33    82.82    83.82    75.57    71.57    61.10    65.37    62.32    65.57    69.84    71.47    65.96    67.72    63.00    63.00    35.83    55.83    71.80    68.80 GPT-4o    89.15    88.29    90.75    89.66    91.33    87.08    84.57    84.57    96.32    97.32    88.55    88.35    87.07    89.44    87.22    83.96    85.59    83.21    90.00    93.33    89.83    86.83    89.73    93.46

### v2

@@ 1 Introduction
Text-to-Image (T2I) generation has emerged as a cornerstone of multimodal AI, enabling the translation of abstract textual concepts into detailed visual content, advancing applications from digital art to scientific visualization. Recent T2I models can be categorized into two main paradigms. Diffusion-based methods—exemplified by Stable Diffusion [1, 2], PixArt [3, 4, 5], Playground [6, 7], FLUX [8], SANA [9, 10, 11], and others[12, 13, 14, 15]—leverage U-Net or Diffusion-Transformer backbones to iteratively denoise Gaussian noises into photorealistic images, delivering superior visual fidelity and diversity. On the other hand, autoregressive (AR) approaches such as LlamaGen [16], Janus [17, 18], Infinity/VAR [19, 20] and other influential open-source efforts [21] treat images as token sequences, employing next-token prediction or scale-progressive generation to synthesize images. Recent works [22, 23, 24, 25] follow the same AR tokenization paradigm but additionally apply reinforcement-learning optimization to further enhance generation quality and controllability. Very recently, commercial T2I models such as GPT-4o [26], Imagen 3 [27] and MidJourney v7 [28] have propelled T2I to a new level. In particular, GPT-4o demonstrates powerful instruction-following capability, which can understand highly intricate prompts and return visually precise, stylistically coherent images in a single conversational loop.
@@ 4.1 Performance of T2I Models on TIIF-Bench
Model    Overall    Basic Following    Advanced Following    Designer          Avg    Attribute    Relation    Reasoning    Avg    Attribute +Relation    Attribute +Reasoning    Relation +Reasoning    Style    Text    Real World    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long    short    long Diffusion based Open-Source Models FLUX.1 dev    71.09    71.78    83.12    78.65    87.05    83.17    87.25    80.39    75.01    72.39    65.79    68.54    67.07    73.69    73.84    73.34    69.09    71.59    66.67    66.67    43.83    52.83    70.72    71.47 SD XL    54.96    42.13    65.72    53.28    59.33    50.83    77.57    62.57    60.32    46.57    49.73    36.22    47.82    35.57    56.22    45.34    52.59    36.09    73.33    60.00    16.83    0.83    50.92    41.59 SD 3    67.46    66.09    78.32    77.75    83.33    79.83    82.07    78.82    71.07    74.07    61.46    59.56    61.07    64.07    68.84    70.34    50.96    57.84    66.67    76.67    59.83    20.83    63.23    67.34 SD 3.5 L    71.15    66.96    78.34    79.56    79.50    76.50    80.96    83.21    72.46    78.71    67.67    61.18    66.46    61.89    73.53    74.15    60.03    61.53    73.33    63.33    70.52    42.52    64.43    66.39 SD3.5 M    70.17    66.19    80.20    75.20    84.50    77.00    78.90    77.51    77.21    71.08    62.33    66.49    66.54    77.51    57.92    62.32    61.46    63.53    80.00    73.33    53.39    28.51    71.64    64.93 FlowGRPO    69.90    69.67    80.15    79.90    89.00    85.00    78.80    78.50    72.64    76.21    63.28    67.33    70.83    72.37    58.37    65.66    59.50    65.85    73.33    66.67    62.44    52.94    64.18    63.81 SANA Sprint    63.68    58.50    76.58    71.00    75.33    71.33    81.82    72.07    72.57    69.57    57.67    51.80    55.32    54.94    68.46    66.72    62.59    63.46    80.00    60.00    8.82    5.83    66.96    58.01 SANA 1.5    67.15    65.73    79.66    77.08    79.83    77.83    85.57    83.57    73.57    69.82    61.50    60.67    65.32    56.57    69.96    73.09    62.96    65.84    80.00    80.00    17.83    15.83    71.07    68.83 Playground v2    45.64    52.78    59.83    69.58    51.33    66.33    70.57    76.07    57.57    66.32    38.43    44.75    41.57    45.57    48.96    59.97    41.72    54.84    53.33    60.00    0.00    0.83    45.32    46.44 Playground v2.5    47.73    54.82    63.08    68.08    57.83    73.83    71.82    77.32    59.57    53.07    40.73    48.17    39.70    45.82    49.59    64.22    44.22    46.72    60.00    80.00    0.00    4.83    47.19    47.56 PixArt-delta    41.01    48.24    53.83    59.25    46.33    52.83    62.07    71.32    53.07    53.57    34.60    42.77    32.44    37.44    53.59    56.59    36.96    49.46    46.67    73.33    0.00    0.00    38.23    40.10 PixArt-alpha    44.37    50.50    55.50    61.00    52.33    56.33    63.82    74.07    50.32    52.57    38.71    44.90    37.82    41.32    58.84    52.46    40.22    47.09    50.00    76.67    0.00    0.83    45.70    53.16 PixArt-sigma    62.00    58.12    70.66    75.25    69.33    78.83    75.07    77.32    67.57    69.57    57.65    49.50    65.20    56.57    66.96    61.72    66.59    54.59    83.33    70.00    1.83    1.83    62.11    52.41 LUMINA-Next    50.93    52.46    64.58    66.08    56.83    59.33    67.57    71.82    69.32    67.07    44.75    45.63    51.44    43.20    51.09    59.72    44.72    54.46    70.00    66.67    0.00    0.83    47.56    49.05 Hunyuan-DiT    51.38    53.28    69.33    69.00    65.83    69.83    78.07    73.82    64.07    63.32    42.62    45.45    50.20    41.57    59.22    61.84    47.84    51.09    56.67    73.33    0.00    0.83    40.10    44.20 AR based Open-Source Models Llamagen    41.67    38.22    53.00    50.00    48.33    42.33    59.57    60.32    51.07    47.32    35.89    32.61    38.82    31.57    40.84    47.22    49.59    46.22    46.67    33.33    0.00    0.00    39.73    35.62 LightGen    53.22    43.41    66.58    47.91    55.83    47.33    74.82    45.82    69.07    50.57    46.74    41.53    62.44    40.82    61.71    50.47    50.34    45.34    53.33    53.33    0.00    6.83    50.92    50.55 Show-o    59.72    58.86    73.08    75.83    74.83    79.83    78.82    78.32    65.57    69.32    53.67    50.38    60.95    56.82    68.59    68.96    66.46    56.22    63.33    66.67    3.83    2.83    55.02    50.92 Infinity    62.07    62.32    73.08    75.41    74.33    76.83    72.82    77.57    72.07    71.82    56.64    54.98    60.44    55.57    74.22    64.71    60.22    59.71    80.00    73.33    10.83    23.83    54.28    56.89 JanusPro    66.50    65.02    79.33    78.25    79.33    82.33    78.32    73.32    80.32    79.07    59.71    58.82    66.07    56.20    70.46    70.84    67.22    59.97    60.00    70.00    28.83    33.83    65.84    60.25 T2I-R1    68.59    67.19    82.90    81.63    86.50    83.00    83.47    79.43    78.73    82.46    69.05    68.00    71.64    69.47    72.43    69.95    69.40    70.40    60.00    63.33    27.60    26.24    67.54    60.45 Closed-Source Models DALL-E 3    74.96    70.81    78.72    78.50    79.50    79.83    80.82    78.82    75.82    76.82    73.39    67.27    73.45    67.20    72.01    71.34    63.59    60.72    89.66    86.67    66.83    54.83    72.93    60.99 MidJourney v6    70.78    67.70    76.00    69.08    77.83    69.33    81.32    73.07    68.82    64.82    68.54    67.62    57.82    61.95    69.84    63.96    57.46    60.34    83.33    73.33    75.83    73.83    65.10    68.46 MidJourney v7    68.74    65.69    77.41    76.00    77.58    81.83    82.07    76.82    72.57    69.32    64.66    60.53    67.20    62.70    81.22    71.59    60.72    64.59    83.33    80.00    24.83    20.83    68.83    63.61 FLUX.1 Pro    67.32    69.89    79.08    78.91    78.83    81.33    82.82    83.82    75.57    71.57    61.10    65.37    62.32    65.57    69.84    71.47    65.96    67.72    63.00    63.00    35.83    55.83    71.80    68.80 GPT-4o    89.15    88.29    90.75    89.66    91.33    87.08    84.57    84.57    96.32    97.32    88.55    88.35    87.07    89.44    87.22    83.96    85.59    83.21    90.00    93.33    89.83    86.83    89.73    93.46